OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. sin 2x + sin x - 2

OpenStudy (souvik):

is it sin2x or sin^2x?

OpenStudy (anonymous):

sin^2x

OpenStudy (souvik):

let x=sin^2x so x^2+x-2.........now factor it..

OpenStudy (anonymous):

should i use the quadratic formula? @souvik

OpenStudy (souvik):

you may do this but u can factor it easily..

OpenStudy (anonymous):

how?

OpenStudy (souvik):

do u know how to factor?

OpenStudy (anonymous):

yeah

OpenStudy (souvik):

come on! try to o it...

OpenStudy (anonymous):

i know how to use the quadratic formula but if there's an easier way i rather do it that way.

OpenStudy (souvik):

okay..did u find x?

OpenStudy (anonymous):

how am i suppose to find x?

OpenStudy (anonymous):

ok so doing the quadratic formula i got\[-1+- \sqrt{7}/2\]

OpenStudy (souvik):

\[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] here a=1 ,b=1 and c=-2

OpenStudy (anonymous):

yeah i did all of that and i got \[-1\pm \sqrt{7}/2\]

OpenStudy (souvik):

no...4ac=4*1*(-2)=(-8) so b^2-4ac=1-(-8)=9

OpenStudy (anonymous):

ok so when i do that i get -1 and -2

OpenStudy (anonymous):

i mean positive 1

OpenStudy (souvik):

right.. So sinx=1...as\[-1\le \sin x \le1\]

OpenStudy (anonymous):

so thats the answer ? sinx=1

OpenStudy (souvik):

sinx=1 or x=90 degree...!

OpenStudy (anonymous):

alright, thanks

OpenStudy (souvik):

welcome..